OpenStudy (anonymous):

Check for the GCF then factor completely by grouping 2x^3 +4x^2 +10x + 20

OpenStudy (anonymous):

gcf = 2 factor by grouping = 2x^2(x + 2) + 10(x+ 2) = (2x^2 + 10)(x+2) = 2(x^2 + 5)(x+ 2)

OpenStudy (anonymous):

oh thanks :) do you do x^3 - 2x^2 + 4^x - 8 the same way??

OpenStudy (anonymous):

yes factor out x^2 for ist two terms then then 4x for second 2

OpenStudy (anonymous):

so it would be (x-2) (x^2 + 4) ?